Mr John wants to load is is house hold on a cargo truck with the help of ramp if the trucks height it is 15m from the ground and the distance from the truck to foot of ramp is 8m find the length of ramp used to load the household in the truck.


1
Expert's answer
2021-05-21T11:07:09-0400

Solution:

Let AB be the truck's height from the ground, AB = 15 m,

And AC be the length of ramp,

And BC be the distance between the foot of truck and foot of the ramp, BC = 8 m.




Triangle ABC is a right-angled triangle, right-angled at B.

Apply Pythagoras theorem,

"AB^2+BC^2=AC^2
\\ \Rightarrow15^2+8^2 =AC^2
\\ \Rightarrow 289=AC^2
\\ \Rightarrow AC=17"

Thus, the length of the ramo is 17 m.